### v1

# Pancharatnam phase as an entanglement witness for quantum gravity in dual Stern-Gerlach interferometers

Abstract
Entanglement plays a central role in fundamental tests and practical applications of quantum mechanics. Because entanglement is a feature unique to quantum systems, its observation provides evidence of the quantum nature on the system. Hence, if gravity can generate entanglement between quantum superpositions, this indicates that quantum amplitudes are field sources and gravity is a quantum phenomenon. I study the dual spin-one-half Stern-Gerlach interferometers proposed to test the quantum nature of gravity and show that the Pancharatnam phase displays features that provide a qualitative distinction between semiclassical and quantum signatures of gravity. The semiclassical evolution of the system is equivalent to that of a single interferometer (a two-level system) in an external field. In this case, a phase jump was observed, as expected from the geodesic rule, which dictates the noncyclic evolution of the superposition in the Bloch sphere. In contrast, in the quantum case, the two interferometers are coupled (two two-level systems) and I find that the phase is continuous.
Introduction: One of the distinctive features of quantum mechanical interactions is their ability to create entanglement between systems. Entangled systems are inherently nonlocal, which fundamentally distinguishes them from their local classical counterparts. However, classical interactions cannot entangle different particles. The expression classical entanglement found in the literature [1, 2] refers to the impossibility of writing a single-particle wave function with different degrees of freedom as a product state. However, even if the single-particle wavefunction cannot be factorized into a product state of these degrees of freedom, physics remains local.
Attempts to integrate general relativity (GR), albeit linearized GR, in the quantum mechanical framework started in the early years of quantum mechanics [3, 4, 5]. However, these efforts aimed at incorporating the geometric formalism of GR into quantum mechanics have not reached completion. A different approach, which is more intuitive and closer to experiments, was suggested by Feynman. One of the experiments he devised was to use the Stern-Gerlach (SG) interferometer to detect a possible quantum signature of gravity [6]. The essential question is whether the quantum amplitudes in the SG apparatus can be sources of gravitational fields. Following this earlier suggestion of laboratory tests of quantum gravity, Bose et al. [7] proposed testing gravity as a quantum phenomenon using entangled dual spin-one-half Stern-Gerlach (SG) interferometers [8, 9, 10]. Another tabletop experiment was proposed in Ref.[11] at the same time. While these proposed experiments still have to overcome technical challenges, they have the merit of defining the quantum gravity problem as a measurable tabletop laboratory experiment. Furthermore, a recent realization of complete SG in cold atom systems has added impetus to the possibility of testing the quantum nature of GR using SG interferometers [12].
Recent matter-wave experiments with external magnetic gradients have shown that a single SG interferometer can generate a noncyclic geometric phase [13]. The geometric phase [14] arising from noncyclic transformations was first discussed theoretically by Samuel and Bhandari [15]. The specific case of a two-level system was extensively explored in subsequent work by Bhandari, who pointed out the occurrence of acute phase jumps during evolution [16, 17]. When the system evolves from point A to point B in the Bloch sphere, the accumulated phase is given by Pancharatnam’s theorem [18]. It is is equal to minus one half of the area enclosed by the trajectory AB and the shortest geodesic joining A and B. The phase jumps result from a sudden change of geodesic when the system passes across specific points in the Bloch sphere. These phase jumps, called $SU(2)$ phase jumps by Bhandari, have been observed experimentally in various systems  [13, 19, 20, 21, 22, 23, 24].
There are two alternatives for interpreting the signal from dual SG interferometers. First, gravity is semiclassical, and the two SG interferometers evolve separately as two spin-one-half subsystems, each under the local gravitational field of the other. This situation is similar to an SG interferometer in an external field as in Ref. [13, 25, 26, 27, 28], the gravitational field playing the role of the magnetic field. For appropriate input parameters, the Pancharatnam phase will display jumps [13] since the system is described by the $SU(2)$ symmetry as in Ref. [16]. Second, gravity is quantized, the two spin-one-half systems are entangled, and the system should be described by a larger subgroup of $SU(4)$, presumably $SO(5)$ because the two two-level systems are identical [29]. As I will show in the following, the Pancharatnam phase is radically different in the two situations, enabling the possibility of clearly distinguishing the quantum behavior of gravity from the semiclassical.
Figure 1: Sketch of the experiment proposed in Ref.  [7]: two nano-particles with equal masses $m$ and spins one-half are released from a trap at $t=0$. After a short time, an equal-weight superposition of each mass was created by an RF pulse of duration $\Delta t_{RF}$. The superposition of the nanoparticles can evolve under mutual gravity. They were then recombined by another RF pulse. The trap release time and $\Delta t_{RF}$ (grey area) are negligible compared to the free-fall time of the two superpositions. Here, I illustrate two possibilities. (A) Quantized gravity: quantum amplitudes are sources of field, and each superposition is sensitive to the gravitational field created by each component of the other superposition (note that the self-gravity term, which in principle should be included, does not induce a phase difference); (B) semiclassical gravity: the quantum amplitudes are not direct sources of gravitational field, and each superposition evolves under a single field created by the effective mass of the other superposition (green oval). In both cases, the interactions are indicated by dotted orange lines.
Dual Stern-Gerlach interferometers: Sketch of the dual Stern-Gerlach interferometers proposed in Ref.  [7] is shown in Fig. 1. Two spin one-half nanoparticles of identical masses, $m$, are released from a trap at $t=0$; then, they are placed in superpositions of equal amplitude by an RF pulse of duration $\Delta t_{RF}$ (grey area), which is small compared to the interferometer time $T$ and will thus be neglected in the calculation of the phase accumulated during evolution. During $T$, the two superpositions evolve under their mutual gravitational fields. If gravity is quantum, as shown in Fig. 1A, each amplitude in the superposition of a first interferometer is a source of a gravitational field. Hence, each branch of the second interferometer is influenced by the fields from the sources at different locations. This leads to entanglement between the two interferometers [7, 30]. However, if I assume that gravity is semiclassical, each superposition creates an effective semiclassical field under which the other field evolves. In the illustration in Fig. 1B the left superposition sees only a field generated by the average mass of the two branches of the right superposition (green oval), and vice versa. The latter situation is similar to the usual studies of a spin-one-half SG interferometer in a magnetic field [13].
Pancharatnam phase: I apply the quantum kinematic theory of Ref. [31] to compute the Pancharatnam phase, assuming the semiclassical or quantum behavior of the dual SG interferometer system. When the system wavefunction evolves from 00 to $t$ the Pancharatnam phase can be obtained from the Bargmann invariant which is given by,
$\Phi(t)=Arg\left[\langle\Psi(0)|\Psi(t)\rangle\right],$ (1)
where $|\Psi(0)\rangle$ and $|\Psi(t)\rangle$ are obtained as in Ref.[7],
$|\Psi(0)\rangle=\frac{1}{2}\left(|\uparrow\uparrow\rangle+|\uparrow\downarrow\rangle+|\downarrow\uparrow\rangle+|\downarrow\downarrow\rangle\right),$ (2)
$|\Psi(0)\rangle=\frac{1}{2}\left(|\uparrow\uparrow\rangle+e^{i\phi_{1}}|\uparrow\downarrow\rangle+e^{i\phi_{2}}|\downarrow\uparrow\rangle+|\downarrow\downarrow\rangle\right).$ (3)
For the quantum case I obtain $\Phi_{Q}$,
$\Phi_{Q}=\arctan\frac{\sin\frac{\phi_{1}+\phi_{2}}{2}\cos\frac{\phi_{1}-\phi_{2}}{2}}{1+\cos\frac{\phi_{1}+\phi_{2}}{2}\cos\frac{\phi_{1}-\phi_{2}}{2}},$ (4)
where,
$\phi_{1}=\frac{\alpha\Delta x}{d(d-\Delta x)},\,\,\phi_{2}=-\frac{\alpha\Delta x}{d(d+\Delta x)}$ (5)
$\alpha=\frac{Gm_{0}^{2}T}{\hbar},\,m=2m_{0}.$ (6)
For the semiclassical case, the calculation of the Pancharatnam phase $\Phi_{c}$ is identical to that of a single interferometer, as in Ref. [13]. I assume that the gravitational field on one interferometer is created by a single mass density on the second interferometer, which is equal to the average mass density of the left and right branches. For equal weights on the branches, this mass is located the center of the interferometer as illustrated in Fig. 1B, and at distances $d-\Delta x/2$ and $d+\Delta x/2$ from the left and right branches of the first interferometer, respectively. I obtain,
$\Phi_{C}=\arctan\frac{\sin\phi}{1+\cos\phi}=\arctan\tan\frac{\phi}{2},$ (7)
where,
$\phi=\frac{\alpha\Delta x}{(d-\frac{\Delta x}{2})(d+\frac{\Delta x}{2})}.$ (8)
Figure 2: The Pancharatnam phase in the dual Stern-Gerlach interferometers experiment proposed in Ref. [7]. The black dashed line corresponds to the semiclassical configuration discussed in Fig. 1A. The continuous blue line corresponds to the quantum configuration of Fig. 1B.
Results: In Fig. 2, I display $\Phi_{C}$ and $\Phi_{Q}$ as functions of the interferometer time. I chose parameters similar to those proposed in Ref. [7]: $m_{0}=5\times 10^{-14}kg,\,d=450\,\mu m,\,\Delta x=250\,\mu m,\,T=1.5\,s$. $\Phi_{C}$ exhibits a sharp phase jump near $T=0.75\,s$. $m_{0}$ was chosen to achieve sufficiently large phase differences at $T=0.75$ s, $\phi_{1}=-0.95$ rad, and $\phi_{2}=6.26$ rad. In semiclassical gravity, $\Phi_{C}$ displays an acute phase jump of magnitude $\pi$, similar to conventional single-SG interferometers under magnetic gradients. Here, the gradients are provided by the effective gravitational field of the superposition (green oval) from one interferometer to the other. These phase jumps can be detected with high accuracy and can be used for precision amplification in metrology [32]. However, when the quantum amplitudes in each branch of the interferometer are considered as sources of the gravitational field, the sharp singularity disappears, as seen in the plot of $\Phi_{Q}$ in Fig. 2. The phase singularity displayed in the semiclassical case becomes an inflection point.
In Fig. 3, I show the visibility as a function of the interferometer time $T$. In the semiclassical system [13] which is dictated by the geodesic rule, the visibility vanishes at the singularity point where the phase $\Phi_{C}$ jumps. This process is illustrated in Fig. 3 where the visibility vanishes near $T=0.75$ s at the position of the singularity. In contrast, in the absence of a singularity, the visibility is not expected to go to zero, it displays a non-zero minimum near $T=0.75$ s.
Figure 3: The visibility in the dual Stern-Gerlach interferometers experiment proposed in Ref. [7]. The black dashed line corresponds to the semiclassical configuration discussed in Fig. 1A. The continuous blue line corresponds to the quantum configuration of Fig. 1B.
Outlook: The Pancharatnam phase provides a means to qualitatively distinguish semiclassical gravity from quantum gravity. The qualitative distinction between these two possible mechanisms for gravity stems from the inherent difference in the symmetry of quantum and semiclassical systems. These are related to two different subgroups of the $SU(4)$ group of the general four-level system. Practical interferometers for gravity tests are likely years away from realization. Here, for the purpose of demonstration, I took $m\sim 10^{-13}$   kg, which yielded sufficient phase accumulation in approximately 1 s to reach the point where the jump occurred. However, these large masses are currently out of reach. As $\Delta\phi\sim m^{2}/d^{2}$, smaller masses can be used by reducing $d$ and achieve large enough phase differences. However, at small values of $d$, the Casimir-Polder (CP) interaction becomes dominant [33]. The CP interaction can be reduced by electromagnetic screening [33]. It can also be reduced by coating the nanoparticles with a near-unity refractive index material. Indices of refraction as low as $n=1.02$ have been reported for silica and $Al_{2}O_{3}$ films [34]. As the CP interaction between two spheres of radius $R$ separated by $d$ is $V_{CP}\sim-\frac{23\hbar c}{4\pi}\frac{R^{6}}{d^{6}}(\frac{\epsilon-1}{\epsilon+2})^{2}$ [33], assuming a small nanodiamond core within a large coating with a low-index material $\epsilon=1.04$, three orders of magnitude can be obtained from a pure nanodiamond for which $\epsilon=5.1$. In addition to CP interactions, other spurious interactions and noise can inhibit the observation of interferometric signals are still being minimized, such as black-body radiation [35], phonons [35], rotational effects [36], and other sources of noise [37].
The author is grateful to Alexandre Bronstein for helpful exchanges. This work was funded in part by the Israel Science Foundation Grants No. 856/18 and No. 1314/19.